OpenStudy (anonymous):

Find the coordinates for the center and length of radius for this circle. (x + 3)2 + (y - 5)2 = 25 center (,) radius

OpenStudy (anonymous):

What is the equation for a circle?

OpenStudy (anonymous):

(x-h squared) + (y-k squared) = r squared

OpenStudy (anonymous):

right. How can you make a plus out of minuses (Hint: x+3 needs to look like x - something)

OpenStudy (anonymous):

x - 6

OpenStudy (anonymous):

no i have to square 3 so it should be 9

OpenStudy (anonymous):

Remember that you need the 3 in there somewhere. x+3 is the same as x - (-3) isn't it?

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

DO NOT EXPAND!!!! The answer is in front of you.

OpenStudy (anonymous):

is it 9 and 25 and radius 5

OpenStudy (anonymous):

the center is 9 and 25 radius 5

OpenStudy (anonymous):

The radius is 5...But you have to find h and k to find the center. That means that you need to make (x+3) look like (x - (-3)). What's h? Then, look at (y-5). Does is look like (y-k)? If so, what's k?

OpenStudy (anonymous):

The center is at (h, k)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

what's k?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Look at the structure:\[(x-h)^{2} +(y-k)^{2}=r^{2}\]

OpenStudy (anonymous):

The minus sign MUST be there. What's k?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

so basically its the same numbers but you have to make them negative

OpenStudy (anonymous):

Sorry. Just 5. No negative. We needed a negative on the 3 for x because there was a plus sign in between, and we needed a minus sign.

OpenStudy (anonymous):

Center at (h, k) = (-3, 5)

OpenStudy (anonymous):

When working with these types of equations, the structure (signs, powers, placement) is the important thing.

OpenStudy (anonymous):

okay can you help with the next one to see if i understand

OpenStudy (anonymous):

Find the coordinates for the center and length of radius for this circle. (x - 3)2 + (y - 5)2 = 8 center (,) radius

OpenStudy (anonymous):

so the answer would be 3, 5 and 2 square root of 2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i am not sure about 2 square root of 2 because it should equally 8

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

this is alot easier than i thought it would be

OpenStudy (anonymous):

But remember that the radius is the square root of what's on the right. Simplifying that is what gave you the 2root 2

OpenStudy (anonymous):

Yeah, it's not bad.

OpenStudy (anonymous):

Find the coordinates for the center and length of radius for this circle. (x - 4)2 + y2 = 12 on this one what do i do with the y squared would it be -1

OpenStudy (anonymous):

There is no k, so k=0